# Denoising Nearest Neighbor Graph via Continuous CRF for Visual Re-ranking without Fine-tuning

### v1

@@ Abstract
Visual re-ranking using Nearest Neighbor graph (NN graph) has been adapted to yield high retrieval accuracy, since it is beneficial to exploring an high-dimensional manifold and applicable without additional fine-tuning. The quality of visual re-ranking using NN graph, however, is limited to that of connectivity, i.e., edges of the NN graph. Some edges can be misconnected with negative images. This is known as a noisy edge problem, resulting in a degradation of the retrieval quality. To address this, we propose a complementary denoising method based on Continuous Conditional Random Field (C-CRF) that uses a statistical distance of our similarity-based distribution. This method employs the concept of cliques to make the process computationally feasible. We demonstrate the complementarity of our method through its application to three visual re-ranking methods, observing quality boosts in landmark retrieval and person re-identification (re-ID).
@@ 1 Introduction
(b) NN graph after reciprocity check
(c) Diffusion-based retrieval of (b)
(d) NN graph denoised by ours
(e) Diffusion-based retrieval of (d)
@@
Figure 1: This figure shows NN graph (affinity) of toy data on manifold. The blue and green colors of nodes indicate different manifolds, while a yellow node and red nodes represent a query point and the retrieval results from a visual re-ranking method of diffusion. The red edges are noisy edges that are connected between different manifolds. Results (e) of our approach (d) are on the correct manifold. Best viewed with zoom-in.
Nearest neighbor search approaches  [1, 2, 3] have used various visual re-ranking techniques  [4, 5, 6] for further improvement of retrieval accuracy. Among them, average query expansion (AQE)  [4] was considered to explore an image manifold on the descriptor space in a simple manner by using a new query averaged over nearest neighbor features at a query time. However, AQE is limited to only information between query and database images.
Contrary to AQE, NN graph based retrieval methods, e.g., diffusion process [7, 8, 6, 9] and graph traversal [10, 11], fully makes use of an affinity matrix between database images as well as information between the query and database images. Diffusion methods, revisited from Donoser et al. [8], constructed the affinity matrix, while considering the query as a part of database images, and calculated solutions of diffusion methods in an iterative manner. However, regarding the query as one of database images requires huge computational costs in online time, because of creating a new affinity matrix whenever new queries are given. Iscen et al. [6] proposed a method that does not have to recreate the matrix even if new queries are given. Yang et al. [12] also made the Iscen’s method more efficient by shifting the diffusion step from online computing to offline one. Recently, Ouyang et al. [13] utilizes self-attention modules to aggregate similarities for improving visual re-ranking outcomes, while requiring additional fine-tuning. Shao et al. [14] highlight visual re-ranking from only top-M results using rich global features to enhance only retrieval precision, not recall. The reliance of this method on the initial ranked list may reduce its robustness across varied or broader retrieval contexts. Our work assumes conditions without fine-tuning and without rich global features, performing re-ranking across the entire database [15].
Yang et al. [16] tackled an issue where NN graph based re-ranking is quite sensitive to noisy data, since all edges, even those containing noise, are used to represent the affinity matrix. Thus, a locally constrained affinity construction was proposed based on edges of satisfying $k$ nearest neighbors ($k$-NN) constraints of images for relaxing noisy edges of negative images. A seminal approach [6] suggested an extension by filtering out noisy edges even from the locally constrained affinity by using a reciprocity check of outgoing and incoming edges of the locally constrained affinity (NN graph).
In this paper, we found that this reciprocity check with NN graph is insufficient to address the noisy edge problem, since the reciprocity check considers only reciprocal relationships between two edges. Since the reciprocity check and $k$-NN constraints do not fully exploit all information from each other, we address the problem with full consideration of relationships between all edges in a clique where we can conservatively reach a consensus (Fig. 1). To realize this, we efficiently employ a fully connected C-CRF, when denoising the noisy NN graph in offline time.
Conditional Random Field (CRF) has been extensively studied [17], and Continuous CRF (C-CRF) [18] was proposed for regression tasks such as depth estimation, image denoising and metric learning for person re-ID [19, 20, 21]. Departing to these applications, our approach targets to refine similarities between each of database images that are robust to noisy edges for improving a diffusion performance in offline time.
Main contributions are summarized as follows:
  • We introduce a C-CRF based refinement method for noisy NN graphs that serves as a pre-processing module to enhance visual re-ranking.
  • Our approach utilizes a statistical distance metric to ensure robust distance measurement and employs cliques for cost-effective denoising.
  • We show that our method significantly enhances existing visual re-ranking methods in landmark retrieval and person re-identification, confirming its role as a complementary tool.
2 Visual re-ranking with NN graph
@@ 3 Approach
Figure 2: An example of clique-based denoising; $k=7$ and $L=10$. (a) shows initial edges of a yellow pivot node, and red edges are noisy edges. (b) shows a clique of the pivot, and all nodes in the clique are fully connected. (c) represents edges denoised by C-CRF, which is constructed from the clique.
@@ 3.1 C-CRF on clique
Our goal is to refine initial similarities of NN graph that are commonly set by top-$k$ similarities of each database image. For computational efficiency and effective noise handling, we utilize the concept of cliques and apply C-CRF into each of cliques. Specifically, we choose a clique, a complete subgraph, where we can have ample information such that we can identify noisy edges. On top of that, using the clique has computational benefits in running CRFs (Sec. 3.2).
@@
Affinity matrix of denoised similarities. Existing diffusion methods [6, 12] employed the reciprocity check over $k$-NN lists for constructing an affinity matrix due to noises and outliers. On the other hand, our denoised similarities of $\bf{y}^{*}$ do not have much of noises and outliers. Thus we can instead use a simple symmetric affinity from the NN graph of our denoised similarities as $a_{pi}=\left(s_{k}(\mathbf{x}_{p}|\mathbf{x}_{i})+s_{k}(\mathbf{x}_{i}|\mathbf{x}_{p})\right)/2$. Afterward, we follow steps of each of diffusion methods with the denoised, symmetric affinity matrix to perform image retrieval.
3.2 Offline complexity
Thanks to the strategic use of cliques, we achieve significant computational efficiency in our C-CRF based denoising process. We outline the computational complexity by detailing each component involved in the process. For the inference stage of a pivot point, we perform three kinds of operations, such as $k$-NN for finding its clique, computing the weight matrix $\bf{W}$ within the clique and calculating the solution for similarity inference. The optimal $k$-NN search takes $\mathcal{O}(k\log N)$ with a whole database size $N$, and it can be accelerated by using some approximated methods [25] if $N$ is sufficiently large. Computing the weight matrix costs $\mathcal{O}(L^{3})$ with the specified clique size $L$ since J-Divergence (Eq. 10) takes $\mathcal{O}(L^{3})$ and lower order terms, such as Euclidean distance, $l_{2}$-normalization and softmax, are disregarded. For similarity inference, it takes $\mathcal{O}(L^{2}t)$ with a specified number of iterations, $t$, for the conjugate gradient method [24]. Note that complexities of these operations are bounded to the fixed clique size $L\ll N$ and the whole complexity for one pivot then becomes a constant time. We perform the same C-CRF refinement for each database image, but this can be parallelized and accelerated by using GPUs and CPUs.
@@ 4 Experiments
In this section, we explore the application of our denoising approach to two distinct but related tasks: landmark retrieval and person re-identification (re-ID). Both tasks benefit from robust image retrieval techniques, making them ideal for demonstrating the versatility of our method.
@@ 4.1 Experimental setup
Datasets and features. We conduct our experiments using challenging datasets: $\mathcal{R}$Oxford and $\mathcal{R}$Paris [26] for landmark retrieval, and DukeMTMC-reID [27] for re-ID. $\mathcal{R}$Oxford and $\mathcal{R}$Paris have three kinds of evaluation protocols, i.e., Easy, Medium, and Hard. Specifically, $\mathcal{R}$Oxford contains $4,993$ images and $70$ queries that represent particular Oxford landmarks and $\mathcal{R}$Paris consists of $6,322$ images and $70$ queries related to particular Paris landmarks. DukeMTMC-reID has 16,522 images with 702 identities for training, 2,228 images with 702 ids for query and 17,661 images for the gallery.
We employ GeM networks [28] of ResNet and VGG versions for landmark retrieval, which are seminal and easy to reproduce in a publicly available author’s implementation [^1] https://github.com/filipradenovic/cnnimageretrieval-pytorch. For re-ID, we leverage ResNet-50 [29], which is publicly available in a repository [^2] https://github.com/layumi/Person_reID_baseline_pytorch, for the baseline network.
Parameter setup. We simply fix C-CRF parameters $\alpha$, $\beta$ to $1$, $0.1$, respectively, for all the following experiments. Parameters $\sigma_{d}$, $\sigma_{r}$ of the Gaussian kernel are set to $0.8$, $2\times 10^{-4}$ for VGG and $0.9$, $3.5\times 10^{-4}$ for ResNet. We empirically choose clique sizes of $1,000$, $500$ and $150$ for $\mathcal{R}$Oxford, $\mathcal{R}$Paris and DukeMTMC-reID datasets, respectively due to the diverse distributions over different datasets. For other baseline methods, we mainly follow the same parameters of each method.
@@ 4.2 Results
Table 1: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Oxford. ResNet and VGG for extracting global features are fine-tuned networks. Improvements without online-time cost are indicated in red, while slight degradations are shown in blue.
@@
Table 2: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Paris. ResNet and VGG for extracting global features are fine-tuned networks. Improvements without online-time cost are indicated in red, while slight degradations are shown in blue.
@@
Table 3: Performance Comparisons against with other re-ranking methods on re-ID setting of DukeMTMC-reID. Rank 1 means a cumulated accuracy at rank-1. Improvements without online-time cost are indicated in red.
Method Rank 1 mAP
NN-Search 64.6 43.6
k-reciprocal 69.3 60.7
EGT 64.5 47.5
+ Ours 64.6(+0.1) 48.2(+0.7)
Online diffusion 68.9 54.9
+ Ours 70.7(+1.8) 62.4(+7.5)
Offline diffusion 68.6 55.0
+ Ours 70.8(+2.2) 62.2(+7.2)
Complementarity Analysis. For validating the denoising quality, we exploit three re-ranking methods, i.e., online diffusion [6], offline diffusion [12], and Explore-Exploit Graph Traversal (EGT) [10]. Tables 1, 2, 3 show performance changes when our method is integrated with other re-ranking methods. Other tested method include Average Query Expansion (AQE) [4] on $R$Oxford and $R$Paris datasets, and k-reciprocal [30] for re-ID setting. In the comprehensive settings analyzed, our method predominantly demonstrated enhancements, with degradation occurring in only two cases, highlighting the high quality of our denoising approach. Overall, the average improvements with our method are 4.6 mAP for $\mathcal{R}$Oxford, 1.7 mAP for $\mathcal{R}$Paris and 3.3 mAP for DukeMTMC-reID. These results are mainly acquired thanks to our C-CRF based denoising.
Table 4: Ablation study of each term constituting our weight function with the Medium protocol, which can represent an overall quality, of each dataset. “w/ ED” and “w/ SD” represent to use the Euclidean distance of CNN-based features and the statistical distance of our similarity-based descriptors, respectively within our C-CRF. Baseline (top row in each feature category w/o any marks) indicates to use only diffusion without C-CRF. The best results are highlighted in bold.
w/ ED w/ SD Offline diffusion Online diffusion
$\mathcal{R}$Oxford $\mathcal{R}$Paris $\mathcal{R}$Oxford $\mathcal{R}$Paris
ResNet 69.9 87.9 67.0 88.5
75.5 89.2 71.4 89.8
75.9 89.3 73.0 91.0
76.1 89.3 73.7 91.0
VGG 70.4 84.1 67.4 85.9
70.9 84.9 67.8 86.2
72.7 85.3 71.0 88.9
73.1 85.4 71.4 89.1
Ablation study. We conduct ablation experiments for each term of the weight function of the C-CRF on both of online diffusion [6] and offline diffusion [12]. The experiments are conducted with the Medium protocol of $R$Oxford and $R$Paris, since the Medium protocol consisting of both of Easy and Hard images that can represent an overall quality of retrieval results. Table 4 displays performances with different weight terms on each network for feature extraction. As shown in the ablation study, we see that using both terms of the Euclidean Distance (ED) of CNN-based features and the Statistical Distance (SD) of similarity-based distributions for the weight function yields the best performance. We then use our C-CRF with ED and SD in the rest of experiments.
@@
Figure 3: mAP graphs on varying $k$. The mAP results are measured on the Medium protocol that reflects the overall performance.
@@
Figure 4: mAP graphs on varying sizes of the clique. The mAP results are measured on the Medium protocol.
(a) Pivot image
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
0ptRefer to caption
[…]
4.3 Qualitative results
[…]
(a) Pivot image
0ptRefer to caption
(b) G-ED:0.36
G-SD:0.65
0ptRefer to caption
(c) G-ED:0.34
G-SD:0.68
0ptRefer to caption
(d) G-ED:0.33
G-SD:0.45
0ptRefer to caption
(e) G-ED:0.37
G-SD:0.35
0ptRefer to caption
(f) G-ED:0.35
G-SD:0.37
0ptRefer to caption
(g) G-ED:0.34
G-SD:0.27
Figure 6: Weights (conformities) used for C-CRF refinement between a pivot image and each of its positive and negative images. Red and blue bars denote negative and positive images of the pivot, respectively. Each of G-ED and G-SD is a weight through the Gaussian kernel (Eq. 11) computed by each distance. The precise weights can be helpful while attaining a right consensus through C-CRF. A higher value means more positive image from the pivot image.
@@ 5 Conclusion
In this study, we introduced a C-CRF based refinement method for denoising NN graph, which is crucial for visual re-ranking. By leveraging cliques, this method enhances computational efficiency while ensuring robust distance measurements through statistical distance metrics for C-CRF. Our approach significantly improves performance in landmark retrieval and person re-identification tasks, demonstrating its effectiveness as a pre-processing tool for enhancing visual re-ranking accuracy.

### v2

# Enhancing Visual Re-ranking through Denoising Nearest Neighbor Graph via Continuous CRF

@@ Abstract
Nearest neighbor (NN) graph based visual re-ranking has emerged as a powerful approach for improving retrieval accuracy, offering the advantages of effectively exploring high-dimensional manifolds without requiring additional fine-tuning. However, the effectiveness of NN graph-based re-ranking is fundamentally constrained by the quality of its edge connectivity, as incorrect connections between dissimilar (negative) images frequently occur. This is known as a noisy edge problem, which hinders the re-ranking performance of existing techniques and limits their potential. To remedy this issue, we propose a complementary denoising method based on Continuous Conditional Random Fields (C-CRF) that leverages statistical distances derived from similarity-based distributions. As a pre-processing step for enhancing NN graph-based retrieval, our approach constructs fully connected cliques around each anchor image and employs a novel statistical distance metric to robustly alleviate noisy edges before re-ranking while achieving efficient processing through offline computation. Extensive experimental results demonstrate that our method consistently improves three different NN graph-based re-ranking approaches, yielding significant gains in retrieval accuracy.
Index Terms—  Visual re-ranking, Visual retrieval, Nearest neighbor search
@@ 1 Introduction
(b) NN graph after
reciprocity check
(c) Diffusion-based
retrieval of (b)
(d) NN graph denoised by
our method
(e) Diffusion-based
retrieval of (d)
@@
Fig. 1: This figure shows NN graph (affinity) of toy data on manifold. The blue and green colors of nodes indicate different manifolds, while a yellow node and red nodes represent a query point and the retrieval results from a visual re- ranking method of diffusion. The red edges are noisy edges that are connected between different manifolds. Results (e) of our approach (d) are on the correct manifold.
Visual retrieval systems typically perform nearest neighbor search with visual descriptors; however, they face fundamental challenges when handling non-linear manifold structures in high-dimensional feature spaces [1, 2]. To address this issue, re-ranking process has emerged as a crucial component in visual retrieval systems, refining initial search results to improve retrieval accuracy. By leveraging various contextual information and relationships between images, re-ranking methods have consistently demonstrated their ability to enhance the quality of retrieval results beyond the initial ranking [3, 4, 5].
Visual re-ranking techniques have evolved along several distinct research directions. Geometric verification approaches [4, 6, 7] leverage local descriptors to verify spatial consistency between image pairs, effectively reducing false matches but incurring significant computational overhead. In contrast, query expansion methods [8, 5, 9] utilize global descriptors to explore image manifolds, aggregating nearest neighbor features to refine the original query descriptor for subsequent retrieval processes. Learning-based approaches [10, 11] employ dedicated re-ranking models with neural networks, however, these methods typically require additional fine-tuning or training data. Meanwhile, nearest neighbor graph-based methods [12, 13] have gained prominence for their ability to explore high-dimensional manifolds without requiring additional training. These approaches, particularly diffusion processes [14, 3] and graph traversal methods [15, 16], achieve superior performance by effectively utilizing comprehensive affinity relationships within the image database.
However, the effectiveness of nearest neighbor graph-based re-ranking is fundamentally constrained by the quality of its edge connectivity in the graph structure. Specifically, incorrect connections between dissimilar (negative) images frequently occur in the initial nearest neighbor graph, known as the noisy edge problem. Such noisy edges significantly hinder re-ranking performance by introducing false connectivity paths that propagate incorrect similarity information throughout the manifold structure. While previous works have attempted to address this issue through locally constrained affinity construction [17] and reciprocity checks [14], these solutions are limited as they only consider pairwise relationships between edges. The reciprocity check, in particular, fails to fully exploit collective information across the graph, as it only considers mutual neighbor relationships while ignoring broader neighborhood structures, leading to suboptimal performance in complex manifold scenarios.
To overcome these limitations, we propose a complementary denoising approach based on Continuous Conditional Random Fields (C-CRF) [18] that leverages statistical distances derived from similarity-based distributions. While CRF has been extensively studied for error reduction in various computer vision tasks such as segmentation and depth estimation [19, 20, 21], its potential for addressing the noisy edge problem in visual retrieval remains unexplored. As a pre-processing step for enhancing NN graph-based retrieval, our method constructs fully connected cliques around each anchor image and considers relationships among all edges within these cliques to reach a more robust consensus, effectively addressing the limitations of pairwise verification approaches. Unlike previous applications of CRF, our approach uniquely focuses on refining pairwise similarities between database images through efficient offline processing, making them robust against noisy edges and thereby improving re-ranking performance. Extensive experimental results demonstrate that our method consistently enhances three different nearest neighbor graph-based re-ranking approaches, yielding significant improvements in retrieval accuracy.
In summary, our contributions are as follows:
  • We propose a complementary denoising method based on C-CRF that enhances NN graph-based re-ranking without requiring additional networks or fine-tuning.
  • We design a statistical distance metric derived from similarity-based distributions and employ fully connected cliques for efficient offline graph denoising.
  • Experimental results demonstrate that our method consistently improves three different NN graph-based re-ranking approaches, confirming its effectiveness as an easily integrable complementary tool.
2 Preliminaries
@@ 3 Approach
Fig. 2: An example of clique-based denoising; $k=7$ and $L=10$. (a) shows initial edges of a yellow pivot node, and red edges are noisy edges. (b) shows a clique of the pivot, and all nodes in the clique are fully connected. (c) represents edges denoised by C-CRF, which is constructed from the clique.
@@ 3.1 C-CRF on clique
Our goal is to refine initial similarities of NN graph that are commonly set by top-$k$ similarities of each database image. For computational efficiency and effective noise handling, we utilize the concept of cliques rather than processing the entire graph. Each clique is then treated with C-CRF to optimize performance. Specifically, we choose a clique, a complete subgraph, where we can have ample information such that we can identify noisy edges.
@@
Affinity matrix of denoised similarities. Existing diffusion methods [14, 22] employed the reciprocity check over $k$-NN lists for constructing an affinity matrix due to noises and outliers. On the other hand, our denoised similarities of $\bf{y}^{*}$ do not have much of noises and outliers. Thus we can instead use a simple symmetric affinity from the NN graph of our denoised similarities as $a_{pi}=\left({\bf{y}}_{p,i}^{*}+{\bf{y}}_{i,p}^{*}\right)/2$. Afterward, we follow steps of each of diffusion methods with the denoised, symmetric affinity matrix to perform visual re-ranking.
@@ 4 Experiments
In this section, we explore the application of our denoising approach to landmark retrieval with three visual re-ranking methods.
@@ 4.1 Experimental setup
Datasets and features. We conduct our experiments using challenging datasets: $\mathcal{R}$Oxford and $\mathcal{R}$Paris [26] for landmark retrieval. $\mathcal{R}$Oxford and $\mathcal{R}$Paris have three kinds of evaluation protocols, i.e., Easy, Medium, and Hard. Specifically, $\mathcal{R}$Oxford contains $4,993$ images and $70$ queries that represent particular Oxford landmarks and $\mathcal{R}$Paris consists of $6,322$ images and $70$ queries related to particular Paris landmarks.
We employ GeM networks [9] of ResNet and VGG versions for landmark retrieval, which are easy to reproduce in a publicly available author’s implementation [^1] https://github.com/filipradenovic/cnnimageretrieval-pytorch.
Parameter setup. We simply fix C-CRF parameters $\alpha$, $\beta$ to $1$, $0.1$, respectively, for all the following experiments. Parameters $\sigma_{d}$, $\sigma_{r}$ of the Gaussian kernel are set to $0.8$, $2\times 10^{-4}$ for VGG and $0.9$, $3.5\times 10^{-4}$ for ResNet. We empirically choose clique sizes of $1,000$ and $500$ for $\mathcal{R}$Oxford and $\mathcal{R}$Paris datasets, respectively. For other baseline methods, we mainly follow the same parameters of each method.
@@ 4.2 Results
Table 1: mean Average Precision (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Oxford. ResNet and VGG for extracting global features are fine-tuned networks. Improvements without online-time cost are indicated in red, while slight degradations are shown in blue.
@@
Table 2: Performance (mAP) comparisons against other re-ranking methods using the NN-graph for $\mathcal{R}$Paris. The layout is the same as in Table 1.
@@
Complementarity Analysis. For validating the denoising quality, we exploit three re-ranking methods, i.e., online diffusion [14], offline diffusion [22], and Explore-Exploit Graph Traversal (EGT) [15]. Tables 1, 2 show performance changes when our method is integrated with other re-ranking methods. Other tested method include Average Query Expansion (AQE) [8] on $R$Oxford and $R$Paris datasets. Overall, the average improvements with our method are 4.6 mAP for $\mathcal{R}$Oxford, 1.7 mAP for $\mathcal{R}$Paris. These results are mainly acquired thanks to our C-CRF based denoising.
Table 3: Ablation study of each term constituting our weight function. “w/ ED” and “w/ SD” represent to use the Euclidean distance of CNN-based features and the statistical distance of our similarity-based descriptors, respectively. Baseline indicates to use only diffusion without C-CRF. The best results are highlighted in bold.
w/ ED w/ SD $\mathcal{R}$Oxford $\mathcal{R}$Paris
ResNet 69.9 87.9
75.5 89.2
75.9 89.3
76.1 89.3
VGG 70.4 84.1
70.9 84.9
72.7 85.3
73.1 85.4
Ablation study. We conduct ablation experiments for each term of the weight function of the C-CRF on offline diffusion [22]. The experiments are conducted with the Medium protocol of $R$Oxford and $R$Paris, since the Medium protocol consisting of both of Easy and Hard images that can represent an overall quality of retrieval results. Table 3 displays performances with different weight terms on each network for feature extraction. As shown in the ablation study, we see that using both terms of the Euclidean Distance (ED) of CNN features and the Statistical Distance (SD) of similarity distributions for the weight function yields the best performance.
@@
Fig. 3: mAP graphs on varying $k$. The mAP results are measured on the Medium protocol.
@@
Fig. 4: mAP graphs on varying sizes of the clique. The mAP results are measured on the Medium protocol.
@@ 5 Conclusion
We propose a C-CRF-based denoising approach to enhance NN graph-based re-ranking. Our method leverages fully connected cliques and a novel statistical distance metric to robustly alleviate noisy edges. Experimental results demonstrate consistent improvements over existing re-ranking approaches, confirming the effectiveness of our method as a complementary tool for enhancing visual re-ranking.
Acknowledgements
This work was supported by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant (No. RS-2023-00237965, Recognition, Action and Interaction Algorithms for Open-world Robot Service) and the National Research Foundation of Korea (NRF) grant (No. RS-2023-00208506), both funded by the Korea government (MSIT). Prof. Sung-Eui Yoon is a corresponding author (e-mail: sungeui@kaist.edu).